# Combating Partial Perception Deficit in Autonomous Driving with Multimodal LLM Commonsense

### v1

@@ 5 Driving-Oriented Fine-tuning
To deliver straightforward optimal driving actions to LLM, we propose fine-tuning an LLM for hazard inference and strategy selection to guide subsequent short-term motion planning. The fine-tuning process involves three tasks: (1) Object Inference, which identifies objects within deficit regions using contextual cues from unaffected parts of the image; (2) Motion Inference, which analyzes motion behaviors in deficit regions based on historical perception frames; and (3) Planning Strategy, which decides between “move” or “stop-observe-move” approaches using insights from the previous tasks. For example, the system may choose “move” with the cue “Bicycle, oncoming at a constant speed,” or “stop-observe-move” with the cue “Pedestrian, crossing to the other side of the road.” Since these three tasks depend on different aspects of the image information, as in Fig. 5, we use three Low Rank Adaptation (LoRA) [16] modules to adapt the model to these tasks.
Figure 6: DriveLM-Deficit Dataset.
We process GVQA dataset from DriveLM-Carla [42] for fine-tuning, which is collected using PDM-Lite [4] expert and achieves 100% completion with zero infractions. GVQA dataset comprises 1.6M QA pairs along with sensor data, keyframes, and frame-level QAs covering perception, prediction, and planning. It employs a rule-based annotation pipeline to determine the ego vehicle’s braking status using privileged information about objects, hazards, and scene measurements. Building on DriveLM-GVQA, we construct the DriveLM-Deficit dataset tailored to our specific case. DriveLM-Deficit consists of 53, 895 videos, each composed of five frames sampled every two steps, forming a 5-second clip at 1 fps. It includes deficits related to traffic lights, traffic signs, pedestrians, and bicycles for fine-tuning. The statistics of DriveLM-Deficit is shown in Fig. 6, and more details are in Appendix A.3.
@@ 6 Experiments
Fine-tuning Details. We choose to supervise fine-tune a lightweight pre-trained vision language model “Qwen2-VL-2B-Instruct” [47] with LoRA using DriveLM-Deficit. we only add LoRA to the “q_proj” and “v_proj” attention layers, and we set the rank to 4 for all experiments.
LLM-RCO Implementation. To control the autonomous agent, we define high-level actions with driving behavior (move forward, stop, change lane to left, change lane to right, turn left, turn right) and speed control (constant speed, deceleration, quick deceleration, deceleration to zero, acceleration, quick acceleration), the action tokens are mapped to control parameters in a deterministic manner based on the navigation to the next target point, more details are provided in Appendix A.1. We employ GPT-4o-mini [1] and fine-tuned Qiwen2-VL-2B as LLM-RCO backbone models.
Perception Deficit Scenarios. To assess the performance of autonomous driving systems under challenging conditions, we designed experiments that simulate different scenarios of perception deficit regarding safety-critical objects. Specifically, using YOLO11 [21], we track and mask safety-critical objects in real-time, including traffic lights, stop signs, pedestrians, and bicycles. When these objects are masked out, control transfers to LLM-RCO.
Carla Simulator. Our implementation employs CARLA (version 0.9.10.1) along with the Transfuser (TF) and Interfuser (IF) models, which integrate both image and LiDAR data for end-to-end vehicle control. Transfuser incorporates rule-based stuck detection and employs a force-move command within its control prediction. Similarly, Interfuser features a rule-based force-move mechanism integrated into its safe controller module. In our setup, both force-move mechanisms remain at their original default settings to enable a fair comparison with LLM-RCO. Experiments are conducted on the Longest6 benchmark, which comprises 36 routes averaging 1.5 km in length and featuring high densities of dynamic actors in diverse environmental conditions.
@@
Visual Deficit LLM-RCO HI-Qiwen,STP-GPT RC $94.15\pm 0.33$ $99.10\pm 5.15$ $100$ $100$ $100$ $95.30\pm 4.17$ $100$ $100$
IS 0.12$\pm$0.04 $0.31\pm 0.06$ 0.03$\pm$0.01 $0.23\pm 0.07$ 0.023$\pm$0.006 $0.19\pm 0.03$ $0.09\pm 0.02$ $0.42\pm 0.08$
DS 11.18$\pm$3.87 $30.4\pm 5.66$ 2.59$\pm$0.73 22.65$\pm$6.91 2.34$\pm$0.64 18.71$\pm$4.05 9.13$\pm$2.45 $41.54\pm 8.10$
AS $2.08\pm 0.03$ $1.82\pm 0.06$ $1.96\pm 0.03$ $1.73\pm 0.02$ $1.15\pm 0.02$ $0.92\pm 0.33$ $1.42\pm 0.08$ $0.97\pm 0.03$
@@
Visual Deficit LLM-RCO HI-GPT,STP-GPT RC 58.57$\pm$3.07 $56.11\pm 1.92$ 37.17$\pm$0.45 $43.15\pm 1.22$ $36.20\pm 0.16$ $41.10\pm 1.32$ $32.50\pm 0.42$ $42.92\pm 0.91$
IS 0.32$\pm$0.06 $0.60\pm 0.07$ $0.24\pm 0.03$ 0.78$\pm$0.09 0.28$\pm$0.02 0.42$\pm$0.09 $0.80\pm 0.13$ 0.78$\pm$0.07
DS 18.58$\pm$3.51 $33.43\pm 4.18$ $8.95\pm 0.72$ 33.70$\pm$4.09 9.94$\pm$0.81 18.79$\pm$4.66 26.26$\pm$4. $32.88\pm 3.01$
Visual Deficit LLM-RCO HI-Qiwen,STP-GPT RC 61.02$\pm$4.88 58.44$\pm$2.31 $36.39\pm 0.87$ $43.60\pm 1.27$ 40.12$\pm$0.70 44.81$\pm$1.33 $30.52\pm 0.62$ 46.80$\pm$1.21
IS $0.28\pm 0.04$ $0.58\pm 0.04$ 0.29$\pm$0.03 $0.74\pm 0.08$ $0.28\pm 0.03$ $0.41\pm 0.08$ 0.81$\pm$0.07 $0.76\pm 0.06$
DS $17.13\pm 2.44$ 33.90$\pm$2.34 10.60$\pm$1.17 $32.26\pm 3.51$ 11.03$\pm$1.66 $18.27\pm 4.40$ $25.02\pm 2.94$ 35.60$\pm$4.56
7 Results
7.1 Fine-tuning results on DriveLM-Deficit
In Fig. 7, we evaluate the hazard inference and motion planning accuracy of GPT-4o-mini, Gemini-2-flash [44], Qiwen2-VL, and fine-tuned Qiwen2-VL on the Drive-Deficit dataset. As shown in Fig.7(a), even advanced LLMs struggle to infer possible objects in occluded or missing image regions under driving scenarios. While they excel at identifying traffic lights and stop signs, they perform poorly in rare cases involving pedestrians and bicycles. To address this limitation, we fine-tuned the open-source Qiwen2-VL model for hazard inference and it achieves higher accuracy on both pedestrians and bicycles than GPT-4o-mini and Gemini2-flash. In Fig. 7(b), we evaluate the LLMs’ ability to infer appropriate motion plan based on past camera frames under perception deficits. The results indicate that advanced LLMs often opt to conservative stop even in scenarios where movement is feasible, which is reflected in lower prediction accuracy for “move” motion strategy across all deficit cases.
@@ 7.1 Fine-tuning results on DriveLM-Deficit
Figure 9: Speed control under various perception deficit scenarios.
7.2 System Results
Overall Results. We report the evaluation results of the LLM-RCO against various perception deficits in two experimental settings: without game time limit shown in Table 1 and with 30-min system time limit shown in Table 2. The game time excludes model inference time, and system time accounts for inference time. In Table 1, we report the average speed (AS) to evaluate the stop rate in the AV movement, where a lower AS indicates more frequent stops during travel. The results in both settings reveal that perception deficits significantly impair autonomous driving performance, leading to remarkably lower driving scores than scenarios without deficits. For Transfuser, the average speed steadily increases with perception deficits, indicating that the loss of critical hazard information leads to inattentive driving. In contrast, for the Interfuser, we observe that the loss of traffic light information results in overly conservative driving due to incorrect traffic light state predictions, causing it to remain stationary until forcibly moved by its rule-based safe controller.
When the LLM-RCO framework intervenes in vehicle control of these adverse scenarios, as shown in Table 1, we observe that the driving score shows improvement with a little compromised route completion score with unlimited game time. In stop sign, pedestrian, and motorcycle cases, the ego vehicle makes more stops, reflected by a decrease in average speed. This cautious driving control of LLM-RCO leads to a higher driving score. For traffic light scenarios, LLM-RCO increases the average speed by reducing unnecessary stops, addressing the issue where incorrect traffic light predictions in Interfuser cause the vehicle to remain stuck until forcibly moved. Moreover, considering the inference time cost of LLM, as shown in Table 2, despite the lower infraction score of LLM-RCO in the traffic light deficit scenario, where the longer completed routes increase exposure to potential risks for transfuser, the infraction and driving scores of LLM-RCO present a consistent increase in all other perception deficit scenarios. The results also demonstrate the fine-tuned Qiwen with more accurate hazard inference and motion planning can reduce stops when the game time is unlimited to enable a complete route and trigger more route completion with limited system time.
Speed Control Analysis. In Fig. 9, we analyze the speed control decisions made by LLM-RCO when overriding the autonomous agent in various perception deficit scenarios. The results are reported with Transfuser on route0 of Town01 in longest6. For scenarios with visual deficits of critical objects, the percentage of different types of speed control remains relatively consistent across test scenarios: approximately $50\%$ of actions maintain a constant speed, around $40\%$ involve deceleration, and only about $10\%$ involve acceleration, indicating that LLM-RCO adopts a cautious driving style, prioritizing safety. Additionally, we observe that the percentage of acceleration actions is lowest in the pedestrian masking scenario, in accordance with rigorous human safety.
@@ 7.2 System Results
Short-term Motion Planning Analysis. In Fig. 8, we examine the performance of LLM-RCO with different limits on planning step lengths. Within the LLM-RCO framework, a maximum length limit is applied to short-term motion planning, enabling the LLM to reason through a variable number of actions based on the current driving context. We report the evaluation metric change with and without LLM-RCO intervention in Transfuser. Our results reveal that extending planning steps increases accident exposure and lowers infraction scores. While a moderate increase in the step limit can improve driving scores, infraction scores drop sharply at higher limits, indicating that excessive planning steps ultimately degrade performance. We also observe that LLM-RCO exhibits different control strategies for static objects (Traffic Lights, Stop Signs) versus moving objects (Bicycles, Pedestrians). When facing static objects, LLM-RCO tends to execute more proactive maneuvers, reflected in higher average speeds and greater route completion (see Tables 2 and 3). Moderately enlarging the planning step limit can further boost driving scores by improving route completion. In contrast, for moving objects, LLM-RCO typically makes conservative stops, which lowers average speeds and hinders route completion. Slightly extending the step length limit prompts the agent to make additional stops, thereby increasing driving scores with increasing infraction scores.
Abalation Study. In Table 3, we conduct ablation studies over the LLM-RCO components based on Transfuser and Interfuser: the Safety Constraint Generator (SC), Hazard Inference Module (HI), Short-Term Motion Planner (ST), and Action Condition Verifier (AC). The results demonstrate the effectiveness of each module. We report the ablation results in the scenario of motorcycle deficit. The results illustrate the contribution of each module to overall driving performance and safety. The test demonstrates that full configuration with all modules active achieves the highest infraction and driving scores. Especially, the core modules of LLM-RCO (HI, ST, AV) are especially necessary to enable valid control override.
@@ 8 Conclusions
Focusing on a realistic scenario in which autonomous vehicles need to ensure safe driving and avoid congestion when their perception is partially compromised, in this paper, we present LLM-RCO, a novel vehicle control framework designed to address this challenge. Our framework harnesses the advanced reasoning capabilities and comprehensive commonsense knowledge inherent in multimodal Large Language Models to enhance autonomous vehicle resilience and safety. By inferring potential hazards within the deficit region and generating corresponding planning-ahead vehicle actions following human-like motion strategies, LLM-RCO can override the vehicle control of autonomous driving agents. Experimental validation demonstrates LLM-RCO’s performance to enable proactive movement despite visual data loss, underlining its potential for modern autonomous driving systems.
Supplementary Material
Appendix A Implementation Details
@@ A.2 Safety-Constraints Generation
For each safety constraint, when the trigger condition is satisfied, we perform the corresponding transformation on the corresponding action $\mathcal{A}_{t}$’s manifold ($\textit{throttle}_{t}$,$\textit{brake}_{t}$ or $\textit{steer}_{t}$). We set two hyperparameters $\Delta_{\text{throttle}}$ and $\Delta_{\text{brake}}$ here, as the transformation factors for the throttle and the brake specifically. Thus, we get the final action manifold:
@@
Compositing Eq. (7), (8), and (9), we arrive at the final action $\mathcal{A}_{t}$ that is executed at time $t$:
@@ A.3 DriveLM-Deficit Dataset
To fine-tune an LLM for hazard inference and motion planning under perception deficits, we post-process DriveLM-GVQA to create DriveLM-Deficit. Specifically, we construct videos using 5 consecutive camera frames from DriveLM-GVQA. The hazard detection and braking annotations of DriveLM-GVQA at the final frame are used to derive the hazard motion and planning strategy label, either “move” or “stop-observe-move”. If the last frame indicates braking, we assign the “stop-observe-move” label; otherwise, we set it to “move”. As shown in Fig.10 (a), in DriveLM-GVQA, the detected hazards are used to determine the brake status for question “Does the ego vehicle need to brake? Why?” For example, “the ego vehicle should stop because of the traffic light that is red.” or “The ego vehicle should stop because of the pedestrian that is crossing the road.” Otherwise, the answer is: ”There is no reason for the ego vehicle to brake.” which corresponds to “move” strategy in our case.
To construct DriveLM-Deficit, we utilize YOLOv11 to track and occlude critical objects in the video, considering two occlusion scenarios. First, as shown in Fig. 10 (b), we directly occlude labeled hazards from DriveLM-GVQA, such as a “stop sign.” In this case, the rationale for the “stop-observe-move” label is: ”The ego vehicle should stop because the invisible region may contain a stop sign.” Second, as shown in Fig. 10 (c), we occlude other objects, such as a “pedestrian” near a stop sign. In this scenario, the ego vehicle must still stop due to the stop sign, regardless of any hazards inferred in the occluded region. We show two data examples of “move” or “stop-observe-move” planning strategy with occluded bicycle in the perception in Fig. 11.

### v2

@@ 5 Driving-Oriented Fine-tuning
To deliver straightforward optimal driving actions to LLM, we propose fine-tuning an LLM for hazard inference and strategy selection to guide subsequent short-term motion planning. The fine-tuning process involves three tasks: (1) Object Inference, which identifies objects within deficit regions using contextual cues from unaffected parts of the image; (2) Motion Inference, which analyzes motion behaviors in deficit regions based on historical perception frames; and (3) Planning Strategy, which decides between “move” or “stop-observe-move” approaches using insights from the previous tasks. For example, the system may choose “move” with the cue “Bicycle, oncoming at a constant speed,” or “stop-observe-move” with the cue “Pedestrian, crossing to the other side of the road.” As shown in Figure 5(a), we use three Low Rank Adaptation (LoRA) (hu2022lora) modules to adapt the model to these tasks.
(a) Driving-Oriented Fine-tuning Scheme.
(b) DriveLM-Deficit Dataset.
Figure 5: Fine-tuning on DriveLM-Deficit.
We process GVQA dataset in DriveLM-Carla (sima2023drivelm) for fine-tuning, which is collected using PDM-Lite (Beißwenger2024PdmLite) and achieves 100% completion with zero infractions. GVQA dataset comprises 1.6M QA pairs along with sensor data, keyframes, and frame-level QAs covering perception, prediction, and planning. It employs a rule-based annotation pipeline to determine the ego vehicle’s braking status using privileged information about objects, hazards, and scene measurements. Building on DriveLM-GVQA, we construct the DriveLM-Deficit dataset tailored to our specific case. DriveLM-Deficit consists of 53, 895 videos, each composed of five frames sampled every two steps, forming a 5-second clip at 1 fps. It includes deficits related to traffic lights, traffic signs, pedestrians, and bicycles for fine-tuning. The statistics of DriveLM-Deficit is shown in Figure 5(b), and more details are in Appendix A.3.
@@ 6 Experiments
Fine-tuning Details. We choose to supervise fine-tune a lightweight pre-trained vision language model “Qwen2-VL-2B-Instruct” (Qwen2VL) with LoRA using DriveLM-Deficit. we only add LoRA to the “q_proj” and “v_proj” attention layers, and we set the rank to 4 for all experiments.
LLM-RCO Implementation. To control the autonomous agent, we define high-level actions with driving behavior (move forward, stop, change lane to left, change lane to right, turn left, turn right) and speed control (constant speed, deceleration, quick deceleration, deceleration to zero, acceleration, quick acceleration), the action tokens are mapped to control parameters in a deterministic manner based on the navigation to the next target point, more details are provided in Appendix A.1. We employ GPT-4o-mini (achiam2023gpt) and fine-tuned Qwen2-VL-2B as LLM-RCO backbone models.
Perception Deficit Scenarios. In the experiments, we simulate different perception deficits of safety-critical objects using YOLO11 (yolo11_ultralytics). We track and mask safety-critical objects in real-time, including traffic lights, stop signs, pedestrians, and bicycles. When these objects are masked out, control transfers to LLM-RCO.
Carla Simulator. Our implementation employs CARLA 0.9.10.1 simulator with the Transfuser (TF) and Interfuser (IF) agents, which integrate both image and LiDAR data for end-to-end vehicle control. Transfuser incorporates rule-based stuck detection and employs a force-move command within its control prediction. Similarly, Interfuser features a rule-based force-move mechanism integrated into its safe controller module. In our setup, both force-move mechanisms remain at their original default settings to enable a fair comparison with LLM-RCO. Experiments are conducted on the Longest6 benchmark, which comprises 36 routes averaging 1.5 km in length and featuring high densities of dynamic actors in diverse environmental conditions.
@@
7 Results
Fine-tuning results on DriveLM-Deficit. In Figure 6, we evaluate the hazard inference and motion planning accuracy of GPT-4o-mini, Gemini-2-flash team2023gemini, Qwen2-VL, and fine-tuned Qwen2-VL on the Drive-Deficit dataset. As shown in Figure 6(a), even advanced LLMs struggle to infer possible objects in occluded or missing image regions under driving scenarios. While they excel at identifying traffic lights and stop signs, they perform poorly in rare cases involving pedestrians and bicycles. To address this limitation, we fine-tuned the open-source Qwen2-VL model for hazard inference and it achieves higher accuracy on both pedestrians and bicycles than GPT-4o-mini and Gemini2-flash. In Figure 6(b), we evaluate the LLMs’ ability to infer appropriate motion plan based on past camera frames under perception deficits. The results indicate that advanced LLMs often opt to conservative stop even in scenarios where movement is feasible, which is reflected in lower prediction accuracy for “move” motion strategy across all deficit cases.
@@ 7 Results
Visual Deficit
LLM-RCO
HI-Qwen,STP-GPT
RC $4.15\pm 0.33 99$$9.10\pm 5.15$ 100100 100100 100100 $95.30\pm 4.17$ 100100 100100
[…]
AS $2.08\pm 0.03$ $1.82\pm 0.06$ $1.96\pm 0.03$ $73\pm 0.02 1.1$$15\pm 0.02$ $0.92\pm 0.33$ $1.42\pm 0.08$ $0.97\pm 0.03$
@@
Visual Deficit
LLM-RCO
HI-GPT,STP-GPT
[…]
Visual Deficit
LLM-RCO
HI-Qwen,STP-GPT
[…]
Overall Results. We report the evaluation results of the LLM-RCO against various perception deficits in two experimental settings: without game time limit shown in Table 1 and with 30-min system time limit shown in Table 2. The game time excludes model inference time, and system time accounts for inference time. In Table 1, we report the average speed (AS) to evaluate the stop rate in the AV movement, where a lower AS indicates more frequent stops. The results in both settings reveal that perception deficits significantly impair autonomous driving performance. For Transfuser, the average speed steadily increases with perception deficits, indicating that the loss of critical hazard information leads to inattentive driving. In contrast, for the Interfuser, we observe that the loss of traffic light information results in overly conservative driving due to incorrect traffic light state predictions, causing it to remain stationary until forcibly moved by its rule-based safe controller.
When LLM-RCO intervenes in vehicle control, as shown in Table 1, we observe that the driving score shows improvement with little compromised route completion score with unlimited game time. In stop sign, pedestrian, and motorcycle cases, the ego vehicle makes more stops, reflected by a decrease in average speed. This cautious driving control of LLM-RCO leads to a higher driving score. For traffic light scenarios, LLM-RCO increases the average speed by reducing unnecessary stops, resolving issues where Interfuser’s incorrect traffic light predictions cause vehicles to remain stuck until forcibly moved. Moreover, considering the inference time cost of LLM, as shown in Table 2, despite the lower infraction score of LLM-RCO in the traffic light deficit scenario, where the longer completed routes increase exposure to potential risks for transfuser, the infraction and driving scores of LLM-RCO present a consistent increase in all other perception deficit scenarios. These results demonstrate the fine-tuned Qwen with more accurate hazard inference and motion planning can effectively reduce stops, enabling more proactive movements.
Figure 7: Speed control analysis.
Speed Control Analysis. In Figure 7, we report the percentage of various speed control decisions of LLM-RCO against various perception deficits. The results are from Transfuser on route0 of Town01 in longest6. The result shows that the percentage of different speed control remains relatively consistent across different deficit scenarios: approximately $50\%$ of actions maintain a constant speed, around $40\%$ involve deceleration, and only about $10\%$ involve acceleration, indicating that LLM-RCO adopts a cautious driving style, prioritizing safety. Additionally, we observe that the percentage of acceleration actions is lowest in the pedestrian masking scenario, in accordance with rigorous human safety.
@@
Short-term Motion Planning Analysis. In Figure 8, we examine the performance of LLM-RCO with different limits on planning step lengths. Within the LLM-RCO framework, a maximum length limit is applied to short-term motion planning, enabling the LLM to reason through a variable number of actions based on the current driving context. We report the evaluation metric change with and without LLM-RCO intervention in Transfuser. Our results reveal that extending planning steps increases accident exposure and lowers infraction scores. While a moderate increase in the step limit can improve driving scores, infraction scores drop sharply at higher limits, indicating that excessive planning steps ultimately degrade performance. We also observe that LLM-RCO exhibits different control strategies for static objects (Traffic Lights, Stop Signs) versus moving objects (Bicycles, Pedestrians). When facing static objects, LLM-RCO tends to execute more proactive maneuvers, reflected in higher average speeds and greater route completion (see Tables 1 and  2). Moderately enlarging the planning step limit can further improve driving scores with more route completion. For moving objects, LLM-RCO typically makes conservative stops, which lowers average speeds and hinders route completion. Slightly extending the step length limit prompts the agent to make additional stops, thereby increasing driving scores with increasing infraction scores.
@@
Ablation Study. In Table 3, we conduct ablation studies over the LLM-RCO components based on Transfuser and Interfuser: the Safety Constraint Generator (SC), Hazard Inference Module (HI), Short-Term Motion Planner (ST), and Action Condition Verifier (AC). We report the ablation results in the scenario of motorcycle deficit. The results illustrate the contribution of each module to overall driving performance and safety. The test demonstrates that full configuration with all modules active achieves the highest infraction and driving scores. Especially, the core modules of LLM-RCO (HI, ST, AV) are especially necessary to enable valid control override.
@@ 8 Conclusions
In this paper, we address a new and challenging problem: enabling proactive safe driving under perception deficits with LLM commonsense and reasoning capabilities. As not all perception deficits are catastrophic to autonomous driving safety, this highlights the need for AVs to be risk-averse, but not entirely risk-avoidant. Therefore, we propose a risk-averse vehicle control framework, LLM-RCO, which incorporates several novel features: action-condition pair generation for verifying action safety with human-crafted rules, a plan-ahead short-term motion planning strategy to reduce latency and inference costs, and safety constraints generation to limit vehicle control based on factors like weather, lighting, and traffic density. Experimental validation demonstrates LLM-RCO’s performance to enable proactive movement despite visual data loss, underlining its potential for modern autonomous driving systems.
Appendix A Appendix
@@ A.2 Safety-Constraints Generation
For each safety constraint, when the trigger condition is satisfied, we perform the corresponding transformation on the corresponding action $\mathcal{A}_{t}$’s manifold ($\textit{throttle}_{t}$$\textit{brake}_{t}$ or $\textit{steer}_{t}$). We set two hyperparameters $\Delta_{\text{throttle}}$ and $\Delta_{\text{brake}}$ here, as the transformation factors for the throttle and the brake specifically. Thus, we get the final action manifold:
@@
Compositing Eq. equation 7, equation 8, and equation 9, we arrive at the final action $\mathcal{A}_{t}$ that is executed at time $t$:
@@ A.3 DriveLM-Deficit Dataset
To fine-tune an LLM for hazard inference and motion planning under perception deficits, we post-process DriveLM-GVQA to create DriveLM-Deficit. Specifically, we construct videos using 5 consecutive camera frames from DriveLM-GVQA. The hazard detection and braking annotations of DriveLM-GVQA at the final frame are used to derive the hazard motion and planning strategy label, either “move” or “stop-observe-move”. If the last frame indicates braking, we assign the “stop-observe-move” label; otherwise, we set it to “move”. As shown in Figure 9 (a), in DriveLM-GVQA, the detected hazards are used to determine the brake status for question “Does the ego vehicle need to brake? Why?” For example, “the ego vehicle should stop because of the traffic light that is red.” or “The ego vehicle should stop because of the pedestrian that is crossing the road.” Otherwise, the answer is: ”There is no reason for the ego vehicle to brake.” which corresponds to “move” strategy in our case.
To construct DriveLM-Deficit, we utilize YOLOv11 to track and occlude critical objects in the video, considering two occlusion scenarios. First, as shown in Figure 9 (b), we directly occlude labeled hazards from DriveLM-GVQA, such as a “stop sign.” In this case, the rationale for the “stop-observe-move” label is: ”The ego vehicle should stop because the invisible region may contain a stop sign.” Second, as shown in Figure 9 (c), we occlude other objects, such as a “pedestrian” near a stop sign. In this scenario, the ego vehicle must still stop due to the stop sign, regardless of any hazards inferred in the occluded region. We show two data examples of “move” or “stop-observe-move” planning strategy with occluded bicycle in the perception in Figure 10.